PROBLEMS:

1. The following persons are under the employment of A & B Enterprises:

Employees Monthly Salary Overtime Work

June Matic P 8,000.00 Reg. Holidays – 8 hrs.

Henry Flores 10,000.00 Rest Day – 10 hrs.

Maria Sarmiento 15,000.00 Reg. Day – 10 hrs.

A & B Enterprises is paying its employees on a semi-monthly basis. The membership fees are deducted on the first half of the month payroll, and the withholding tax is deducted on the second half of the month payroll. The employees did not incur any tardiness and absences during the month. They are entitled to overtime pay. The enterprise is using the 314 working days in a year as basis with 8 regular working hours in a day.

Requirements:

1. Compute the withholding faxes of each employee.

2. Compute the overtime pay of each employee.

3. Prepare the payroll sheet for the period March 1 – 15, 2020.



1
Expert's answer
2021-03-08T09:23:08-0500

1.Only 30 days in June. Let the working hours be 22 days, established by law. The norm of hours about 8 of the working week in June is 167 hours

Let overtime be paid 2 times for each hour.

Let the rate per hour be 500, then we will calculate the hours of overtime:

"8000=2\times x\times500"

x=8 Matic P

"10 000=2\times x\times500"

x=10 Henry Flores

"15 000=2\times x\times500"

x=15 Maria Sarmiento

The main hourly work then amounts to

"S=h\times T=167\times500=83500"

2.Let the membership fees be 1000 and the tax be 10%

tax:

"(83500+8000)\times0.1=9150"

"(83500+1000)\times0.1=9350"

"(83500+15000)\times0.1=9850"

3.

We will make up the payroll sheet:

The advance is calculated as 40 percent of the accrued income minus: tax and contributions


 Postscript: Interest, tariff, etc. can be changed and there will be other amounts